# Arbitrarily-Conditioned Multi-Functional Diffusion for Multi-Physics Emulation

### v1

@@ 5.2 Generation Performance
We then evaluated the generation performance of ACM-FD by using it to generate instances of all relevant functions for the Darcy Flow and Convection Diffusion systems. We compared ACM-FD with a popular version of the Variational Auto-Encoder (VAE), known as $\beta$-VAE (Higgins et al., 2022), where $\beta$ controls the regularization strength from the prior. The encoder network was built using a series of convolutional layers, while the decoder network was constructed as the transpose of these layers. For training and hyperparameter tuning of $\beta$-VAE, we used the same training and validation sets as those used by ACM-FD in Section 5.1. The details are provided in Section B. To let $\beta$-VAE train with irregularly-sampled data, we also applied scipy.interpolate.griddata for interpolation.
@@
We used two metrics to evaluate the quality of the generated data. The first is the equation error, which measures how well the generated data adheres to the governing equations of the system. To calculate this error, for each group of sampled source function, parameter function, and/or initial conditions, we ran the numerical solver (the same one used to generate the training data) to solve the governing equation(s), and then computed the relative $L_{2}$ error between the generated solution and the numerical solution. The second metric is the Mean-Relative-Pairwise-Distance (MRPD), which measures the diversity of the generated data and is commonly used in the evaluation of generative models  (Yuan and Kitani, 2020; Barquero et al., 2023; Tian et al., 2024).
(a) Darcy Flow
(b) Convection Diffusion
(c) Torus Fluid
Figure 1: Generated Function Instances By ACM-FD (top row) And By $\beta$-VAE (Second Row). "-sol" Means The Numerical Solution Provided By The Numerical Solvers Given The Other Functions.
Table 2: Relative $L_{2}$ Error For Various Prediction Tasks, With Irregularly Sampled Training And Test Functions. The Results Were Averaged Over Five Runs.
System Task(s) ACM-FD FNO GNOT DON PODDON
D-F $f,u$ to $a$ 1.35e-02 (3.33e-04) 2.79e-02 (2.89e-04) 1.35e-01 (2.18e-04) 3.39e-02 (6.71e-04) 3.59e-02 (8.51e-04)
$a,u$ to $f$ 1.08e-02 (1.11e-02) 2.28e-02 (2.21e-02) 1.00e+00 (1.00e+00) 3.23e-02 (3.26e-02) 4.28e-02 (4.16e-02)
$a,f$ to $u$ 1.33e-02 (8.12e-05) 4.49e-02 (4.79e-04) 9.99e-01 (9.27e-04) 5.34e-02 (3.85e-04) 6.58e-02 (7.26e-04)
$u$ to $a$ 3.82e-02 (8.35e-04) 5.91e-02 (5.68e-04) 1.35e-01 (9.39e-05) 9.42e-02 (3.15e-03) 1.01e-01 (1.02e-03)
$u$ to $f$ 3.77e-02 (4.45e-04) 6.07e-02 (5.58e-04) 9.99e-01 (7.53e-04) 1.01e-01 (3.26e-03) 9.47e-02 (1.34e-03)
C-D $s,u$ to $v$ 2.65e-02 (7.79e-04) 5.55e-02 (6.98e-04) 1.05e-01 (2.19e-03) 4.37e-02 (2.19e-04) 5.91e-02 (9.05e-04)
$v,u$ to $s$ 5.87e-02 (1.29e-03) 9.87e-02 (9.71e-04) 1.46e-01 (2.24e-03) 9.01e-02 (3.01e-03) 1.18e-01 (1.06e-03)
$v,s$ to $u$ 3.45e-02 (3.87e-04) 6.64e-02 (6.63e-05) 2.40e-02 (4.67e-04) 7.59e-02 (2.76e-04) 2.17e-01 (6.16e-04)
$u$ to $v$ 3.11e-02 (2.06e-04) 7.21e-02 (1.64e-03) 1.86e-01 (1.64e-02) 6.03e-02 (1.86e-03) 7.79e-02 (1.43e-03)
$u$ to $s$ 6.34e-02 (8.90e-04) 1.36e-01 (2.24e-03) 2.04e-01 (6.26e-03) 1.22e-01 (3.56e-03) 1.77e-01 (3.13e-03)
T-F $w_{0},w_{5}$ to $w_{1}$ 1.78e-02 (1.47e-03) 4.66e-02 (1.18e-03) 6.36e-02 (1.36e-03) 7.72e-02 (8.98e-04) 8.97e-02 (8.82e-04)
$w_{0},w_{5}$ to $w_{2}$ 1.52e-02 (5.54e-04) 5.10e-02 (8.67e-04) 7.81e-02 (1.10e-03) 9.04e-02 (4.75e-04) 1.05e-01 (1.10e-03)
$w_{0},w_{5}$ to $w_{3}$ 1.67e-02 (1.75e-03) 5.42e-02 (4.75e-04) 8.34e-02 (8.63e-04) 8.76e-02 (4.50e-04) 1.05e-01 (9.14e-04)
$w_{0},w_{5}$ to $w_{4}$ 1.34e-02 (1.47e-03) 4.67e-02 (2.52e-04) 7.06e-02 (4.67e-04) 7.12e-02 (2.48e-04) 8.60e-02 (5.36e-04)
$w_{0},w_{5}$ to $f$ 1.45e-02 (5.65e-04) 5.92e-02 (3.72e-04) 8.56e-02 (2.48e-03) 6.74e-02 (7.74e-04) 9.62e-02 (9.64e-04)
$w_{0},f$ to $w_{1}$ 1.96e-02 (1.97e-03) 4.32e-02 (1.98e-04) 5.89e-02 (7.27e-04) 7.72e-02 (5.28e-04) 8.68e-02 (8.65e-04)
$w_{0},f$ to $w_{2}$ 2.10e-02 (2.17e-03) 4.75e-02 (5.11e-04) 7.35e-02 (1.73e-03) 9.78e-02 (1.26e-03) 1.12e-01 (1.05e-03)
$w_{0},f$ to $w_{3}$ 2.03e-02 (1.50e-03) 5.39e-02 (2.97e-04) 9.73e-02 (1.86e-03) 1.13e-01 (6.47e-04) 1.31e-01 (6.74e-04)
$w_{0},f$ to $w_{4}$ 2.13e-02 (1.17e-03) 6.07e-02 (1.02e-03) 1.03e-01 (2.57e-03) 1.34e-01 (1.19e-03) 1.49e-01 (1.11e-03)
$w_{0},f$ to $w_{5}$ 2.51e-02 (1.39e-03) 6.66e-02 (8.54e-04) 1.10e-01 (5.86e-03) 1.64e-01 (2.67e-03) 1.76e-01 (1.53e-03)
@@
For each system, we generated 1,000 sets of the relevant functions. The average equation error and MRPD are presented in Table 3. As shown, the functions generated by ACM-FD consistently exhibit relatively small equation errors, indicating a strong alignment with the physical laws governing each system. In contrast, the equation errors for the data generated by $\beta$-VAE are much larger, often by an order of magnitude. Furthermore, the data generated by ACM-FD demonstrates a higher MRPD than those from $\beta$-VAE, indicating better diversity among the generated function instances. Overall, these results suggest that ACM-FD produces not only reliable but also diverse multi-physics data.
Next, we investigated the functions generated by each method. Figure 1 illustrates a set of functions randomly generated by ACM-FD and $\beta$-VAE for each system. More examples are provided in Appendix Figure 2. Our method’s generated solutions are highly consistent with the numerical solutions, accurately capturing both the global structures and finer details. In contrast, the solutions produced by $\beta$-VAE often fail to capture local details (see Fig. 1(a) for $u$ and Fig. 1(c) for $w_{1}$) or significantly deviate from the numerical solution in terms of overall shape (see Fig. 1(b) for $u$ and Fig. 1(c) for $w_{3}$).
In addition, other functions generated by $\beta$-VAE, such as $(a,f)$ for Darcy Flow and $(v,s)$ for Convection Diffusion, appear quite rough and do not align well with the corresponding function families (see Appendix Section A). This might be due to that $\beta$-VAE is unable to capture the correlations between function values across different input locations. In contrast, ACM-FD introduces GP noise functions and performs diffusion and denoising within the functional space. As a result, ACM-FD is flexible enough to effectively capture the diverse correlations between function values. Overall, these results confirm the advantages of our method in multi-physics data generation.
@@
Table 3: Equation Error And Diversity Of Generated Data For The Whole System. D-F, C-D, And T-F Are Short For Darcy Flow, Convection Diffusion, And Torus Fluid, Respectively. MRPD Is Short For Mean Relative Pairwise Distance.
System (setting) Task(s) ACM-FD $\beta$-VAE
D-F (Regular) Equation Error 0.0576 0.265
MRPD 1.15 0.932
C-D (Regular) Equation Error 0.114 0.282
MRPD 1.00 0.879
T-F (Regular) Equation Error 0.0273 0.737
MRPD 0.8042 0.524
D-F (Irregular) Equation Error 0.0427 0.32
MRPD 1.52 0.928
C-D (Irregular) Equation Error 0.065 0.285
MRPD 1.49 0.949
T-F (Irregular) Equation Error 0.0221 0.741
MRPD 0.916 0.4907
@@ 6 CONCLUSION
We have presented ACM-FD, a novel probabilistic generative model designed for multi-physics emulation. ACM-FD can perform arbitrarily-conditioned multi-function sampling and has demonstrated potential in several classical systems. In future work, we plan to extend our investigations into more complex multi-physics systems and explore alternative architecture designs to further enhance its capabilities.
@@
A Mulit-Physics Systems Details
@@ A.1 Darcy Flow
where $a({\bf x})$ is the permeability field, $u({\bf x})$ is the fluid pressure, and $f({\bf x})$ is the external source, representing sources or sinks of the fluid. To obtain one instance of the triplet $(a,f,u)$, we sampled $f$ from a Gauss random field: $f\sim\mathcal{N}(0,(-\Delta+25)^{-\frac{15}{2}})$, and $a$ from the exponential of a Gaussian random field: $a=\exp(g)$ where $g\sim\mathcal{N}(0,(-\Delta+16)^{-2})$. We then ran a second-order finite-difference solver to obtain $u$ at a $256\times 256$ mesh. To prepare irregularly-sampled data, we randomly selected 1,200 to 1,500 input locations and their corresponding function values from the mesh. To prepare regularly-sampled data, we extracted each function instance from a $64\times 64$ sub-mesh.
@@ A.2 Convection Diffusion
where $(x,t)\in[-1,1]\times[0,1]$, $u(x,0)=0$, $D$ is the diffusion coefficient and was set to 0.01, $v(x,t)$ is the convection velocity, describing how fast the substance is transported due to the flow, $s(x,t)$ is the source term, representing the external force, and $u(x,t)$ is the quantity of interest, such as the temperature, concentration and density. We employed a parametric form for $v$ and $a$: $v(x,t)=\sum_{n=1}^{3}a_{n}x^{n}+a_{4}t$, and $s(x,t)=\alpha\exp(-\beta(x+t)^{2})+\gamma\cos(\eta\cdot\pi(x-0.1t))$, where all $a_{n}$, $\alpha,\beta,\gamma,\eta$ are sampled from $\text{Uniform}(-1,1)$. We used the MATLAB PDE solver pdepe [^1] https://www.mathworks.com/help/matlab/math/partial-differential-equations.html to obtain the solution for $u$. The spatial-temporal domain was discretized into a $512\times 256$ mesh. To prepare irregularly-sampled data, we randomly selected 1,200 to 1,500 input locations and their corresponding function values from the mesh. To prepare regularly-sampled data, we extracted each function instance from a $64\times 64$ sub-mesh.
A.3 Torus Fluid
@@
where ${\bf x}\in[0,1]^{2}$, $t\in[0,10]$, $\nu=0.001$, ${\bf u}$ is the velocity field that $\nabla\cdot{\bf u}=0$, $w({\bf x},t)$ is the vorticity function, and $f({\bf x})$ is the source function that represents the external forces. We are interested in seven spatial functions: the initial condition $w_{0}$, the source function $f$, and the vortocity fields at time steps $t=2,4,6,8,10$, denoted as $w_{1}$, $w_{2}$, $w_{3}$, $w_{4}$, and $w_{5}$. We employed a parametric form for $w_{0}$ and $f$: $w_{0}({\bf x})=\left[\sin(\alpha_{1}\pi(x_{1}+\beta_{1})),\sin(\alpha_{2}\pi(x_{1}+\beta_{2}))\right]\cdot\mathbf{\Lambda}\cdot\left[\cos(\alpha_{1}\pi(x_{2}+\beta_{1})),\cos(\alpha_{2}\pi(x_{2}+\beta_{2}))\right]^{T}$, where $\alpha_{1},\alpha_{2}\sim\text{Uniform}(0.5,1)$, $\beta_{1},\beta_{2}\sim\text{Uniform}(0,1)$, and each $\left[\mathbf{\Lambda}\right]_{ij}\sim\text{Uniform}(-1,1)$, and $f(x_{1},x_{2})=0.1\left(a\sin(2\pi(x_{1}+x_{2}+c))+b\cos(2\pi(x_{1}+x_{2}+d)\right)$ where $a,b\sim\text{Uniform}(0,2)$ and $c,d\sim\text{Uniform}(0,0.5)$. We used the finite-difference solver as provided in (Li et al., 2020b) to obtain the solution of $\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. The spatial domain was discretized to a $128\times 128$ mesh. Again, to prepare the irregularly-sampled training and test data, we randomly selected 1,200 to 1,500 locations and the corresponding function values from the mesh. To prepare the regularly-sampled data, we extracted the function instances from a $64\times 64$ sub-mesh.
B Hyperparameter Selection

### v2

@@ 5.2 Generation Performance
We then evaluated the generation performance of ACM-FD by using it to generate instances of all relevant functions for the Darcy Flow, Convection Diffusion and Torus Fluid systems. We compared ACM-FD with two methods. The first one is multi-function diffusion (MFD) solely for generating all the functions. We used the same architecture as ACM-FD but trained the model unconditionally, in accordance with the standard DDPM framework (see (9)). The second method is $\beta$-VAE (Higgins et al., 2022), a widely used Variational Auto-Encoder (VAE), where $\beta$ controls the regularization strength from the prior. The encoder network was built using a series of convolutional layers, while the decoder network was constructed as the transpose of these layers. For training and hyperparameter tuning of $\beta$-VAE, we used the same training and validation sets as those used by ACM-FD in Section 5.1. The details are provided in Appendix Section B.
@@
We used two metrics to evaluate the quality of the generated data. The first is the equation error, which measures how well the generated data adheres to the governing equations of the system. To calculate this error, for each group of sampled source function, parameter function, and/or initial conditions, we ran the numerical solver — the same one used to generate the training data — to solve the governing equation(s), and then computed the relative $L_{2}$ error between the generated solution and the numerical solution. The second metric is the Mean-Relative-Pairwise-Distance (MRPD), which measures the diversity of the generated data (Yuan & Kitani, 2020; Barquero et al., 2023; Tian et al., 2024). It is worth noting that since there is no large pretrained inception network tailored to encompass all types of functions across numerous multi-physics systems, the popular FID score (Heusel et al., 2017) is not applicable to our evaluation.
@@
For each system, we generated 1,000 sets of the relevant functions. The average equation error and MRPD are presented in Table 2. As shown, the functions generated by ACM-FD and MFD consistently exhibit relatively small equation errors, indicating strong alignment with the physical laws governing each system. In contrast, the equation errors for functions generated by $\beta$-VAE are significantly larger, often by an order of magnitude. Furthermore, the data generated by ACM-FD and MFD demonstrates a higher MRPD compared to $\beta$-VAE, indicating better diversity among the generated function instances. For the functions generated by ACM-FD and MFD, both the equation error and MRPD are comparable. Specifically, in Darcy Flow (D-F) and Convection Diffusion (C-D), ACM-FD achieves smaller equation errors and higher MRPD, while in Torus Fluid (T-F), the metrics for ACM-FD are slightly worse. These results suggest that ACM-FD, despite being trained for a variety of prediction and conditional generation tasks, still achieves comparable unconditional generation performance to the diffusion model trained purely for unconditional generation. Overall, these findings indicate that ACM-FD is capable of producing multi-physics data that is not only reliable but also diverse. In Appendix Section C, we provide examples of the functions generated by each method, along with detailed discussion and analysis.
@@
Table 2: Equation relative $L_{2}$ error and diversity of generated data for the whole system. MRPD is short for Mean Relative Pairwise Distance.
System Task(s) ACM-FD MFD $\beta$-VAE
D-F Equation Error 0.0576 0.0584 0.265
MRPD 1.15 0.980 0.932
C-D Equation Error 0.114 0.127 0.282
MRPD 1.00 0.971 0.879
T-F Equation Error 0.0273 0.0234 0.737
MRPD 0.8042 0.9537 0.524
Table 3: Relative $L_{2}$ error for function completion. Each function is sampled in one half of the domain, while the other half is completed using different methods.
Dataset Task(s) ACM-FD MFD-Inpaint Interp
D-F $a$ 1.21e-02 7.94e-02 1.04e-01
$f$ 1.23e-02 6.41e-02 6.98e-01
$u$ 1.09e-02 2.71e-02 8.07e-01
C-D $v$ 1.87e-02 4.71e-01 8.30e-01
$s$ 3.39e-02 3.22e-01 6.49e-01
$u$ 1.45e-02 3.47e-02 8.97e-01
5.3 Completion Performance
Third, we evaluated ACM-FD on the task of completing functions in unobserved domains. Specifically, we tested ACM-FD on Darcy Flow (D-F) and Convection Diffusion (C-D) systems. For each function, we provided samples from half of the domain and used ACM-FD to predict the function values in the other half. To benchmark performance, we adapted our multi-functional diffusion (MFD) framework to the DDPM inpainting approach proposed in (Lugmayr et al., 2022), which we refer to as MFD-Inpaint. This method trains MFD unconditionally as in Section 5.2. During the completion process, the observed function values are perturbed using the forward process, and these perturbed values, along with noise for the unobserved function values, are passed into the denoising network. Additionally, we compared ACM-FD with the interpolation strategy, which is commonly used in scientific and engineering domains. We employed scipy.interpolate.griddata for interpolation. We denote this method as Interp. The relative $L_{2}$ error for each method is presented in Table 3. As shown, ACM-FD consistently achieves much lower errors compared to both MFD-Inpaint and Interp. The superior performance of ACM-FD over MFD-Inpaint might be attributed to its explicitly conditional training strategy, which allows ACM-FD to leverage conditioned information more effectively during completion. It is also noteworthy that both ACM-FD and MFD-Inpaint substantially outperform classical interpolation, demonstrating that diffusion-based modeling is far more effective for inferring unknown function values. Completion examples for each method are visualized in Appendix Fig. 6.
Table 4: Empirical Coverage Probability (ECP) under confidence levels $\alpha\in\{0.9,0.95,0.99\}$. Best results per task are in boldface.
Dataset Task Method 0.9 0.95 0.99
C-D $s,u$ to $v$ ACM-FD 0.833 0.880 0.921
Simformer 0.736 0.814 0.871
$v,u$ to $s$ ACM-FD 0.766 0.842 0.913
Simformer 0.683 0.767 0.879
$v,s$ to $u$ ACM-FD 0.939 0.968 0.990
Simformer 0.695 0.771 0.858
$u$ to $v$ ACM-FD 0.821 0.870 0.922
Simformer 0.775 0.850 0.912
$u$ to $s$ ACM-FD 0.920 0.949 0.972
Simformer 0.716 0.773 0.823
D-F $a,u$ to $f$ ACM-FD 0.947 0.974 0.991
Simformer 0.829 0.895 0.950
$a,f$ to $u$ ACM-FD 0.985 0.994 0.998
Simformer 0.922 0.955 0.998
$u$ to $f$ ACM-FD 0.867 0.909 0.952
Simformer 0.918 0.953 0.980
5.4 Uncertainty Quantification
Fourth, we assessed the quality of uncertainty calibration provided by ACM-FD. To this end, we examined the empirical coverage probability (Dodge, 2003): $\text{ECP}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(y_{i}\in C_{\alpha})$, where $y_{i}$ is the ground-truth function value, and $C_{\alpha}$ is the $\alpha$ confidence interval derived from 100 predictive samples generated by our method. We varied $\alpha$ from {90%, 95%, 99%}, and examined our method in eight prediction tasks across the Convection-Diffusion (C-D) and Darcy Flow (D-F) systems. We compared against Simformer. Note that all the other competing methods are deterministic and therefore unable to perform uncertainty quantification. As shown in Table  4, in most cases our method achieves coverage much closer to $\alpha$, showing superior quality in the estimated confidence intervals.
Additionally, we visualized prediction examples along with their uncertainties, measured by predictive standard deviation (std). As shown in Figure 2, more accurate predictions tend to have lower std, i.e., low uncertainty, while regions with larger prediction errors correspond to higher std — providing further qualitative evidence that our uncertainty estimates are well-aligned with prediction quality.
(a) D-F: $u$ to $f$.
(b) D-F: $a$, $u$ to $f$.
(c) D-F: $a$, $f$ to $u$.
(d) C-D: $s$, $u$ to $v$.
(e) C-D: $v$, $u$ to $s$.
Figure 2: Visualizations of ACM-FD predictions.
5.5 Ablation Studies
Finally, we conducted ablation studies to assess the contribution of the key components in ACM-FD, including random masking strategy, masking probability $p$, and Kronecker product-based computation. The results and analysis, as presented in Appendix Section D, confirm the effectiveness of each component.
@@ 6 Conclusion
We have presented ACM-FD, a novel probabilistic generative model designed for multi-physics emulation. ACM-FD can perform arbitrarily-conditioned multi-function sampling and has demonstrated potential in several classical systems. In future work, we plan to extend our investigations into more complex multi-physics systems and explore alternative architecture designs to further extend its capabilities.
Acknowledgements
SZ, DL and ZX acknowledge the support of MURI AFOSR grant FA9550-20-1-0358, NSF CAREER Award IIS-2046295, NSF OAC-2311685, and Margolis Foundation. AN acknowledges the support of the grant AFOSR FA9550-23-1-0749. This work has also leveraged NCSA Delta GPU cluster from NCF Access program under Project CIS240255.
Impact Statement
This paper presents work whose goal is to advance machine learning for multi-physics emulation. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgements
Appendix A Mulit-Physics Systems Details
@@ A.1 Darcy Flow
where $a({\bf x})$ is the permeability field, $u({\bf x})$ is the fluid pressure, and $f({\bf x})$ is the external source, representing sources or sinks of the fluid. To obtain one instance of the triplet $(a,f,u)$, we sampled $f$ from a Gauss random field: $f\sim\mathcal{N}(0,(-\Delta+25)^{-\frac{15}{2}})$, and $a$ from the exponential of a Gaussian random field: $a=\exp(g)$ where $g\sim\mathcal{N}(0,(-\Delta+16)^{-2})$. We then ran a second-order finite-difference solver to obtain $u$ at a $256\times 256$ mesh. We extracted each function instance from a $64\times 64$ sub-mesh.
@@ A.2 Convection Diffusion
where $(x,t)\in[-1,1]\times[0,1]$, $u(x,0)=0$, $D$ is the diffusion coefficient and was set to 0.01, $v(x,t)$ is the convection velocity, describing how fast the substance is transported due to the flow, $s(x,t)$ is the source term, representing the external force, and $u(x,t)$ is the quantity of interest, such as the temperature, concentration and density. We employed a parametric form for $v$ and $a$: $v(x,t)=\sum_{n=1}^{3}a_{n}x^{n}+a_{4}t$, and $s(x,t)=\alpha\exp(-\beta(x+t)^{2})+\gamma\cos(\eta\cdot\pi(x-0.1t))$, where all $a_{n}$, $\alpha,\beta,\gamma,\eta$ are sampled from $\text{Uniform}(-1,1)$. We used the MATLAB PDE solver pdepe [^1] https://www.mathworks.com/help/matlab/math/partial-differential-equations.html to obtain the solution for $u$. The spatial-temporal domain was discretized into a $512\times 256$ mesh. We extracted the function instances of interest from a $64\times 64$ equally-spaced sub-mesh.
A.3 Diffusion Reaction
We used the 2D diffusion-reaction dataset provided from PDEBench (Takamoto et al., 2022). The dataset was simulated from a 2D diffusion-reaction system, which involves an activator function $v_{1}(t,x,y)$ and an inhibitor function $v_{2}(t,x,y)$. These two functions are non-linearly coupled. The governing equation is given as follows.
$\frac{\partial v_{1}}{\partial t}$ $=D_{1}\frac{\partial^{2}v_{1}}{\partial x^{2}}+D_{1}\frac{\partial^{2}v_{1}}{\partial y^{2}}+v_{1}-v_{1}^{3}-k-v_{2},$
$\frac{\partial v_{2}}{\partial t}$ $=D_{2}\frac{\partial^{2}v_{2}}{\partial x^{2}}+D_{2}\frac{\partial^{2}v_{2}}{\partial y^{2}}+v_{1}-v_{2},$ (16)
where $x,y\in(-1,1)$, $t\in(0,5]$, $k=0.005$, and the diffusion coefficients $D_{1}=0.001$ and $D_{2}=0.005$. The initial condition is generated from the standard Gaussian distribution. The numerical simulation process is detailed in (Takamoto et al., 2022). We are interested in four functions: $f_{1}=v_{1}(2.5,x,y)$, $f_{2}=v_{2}(2.5,x,y)$, $u_{1}=v_{1}(5.0,x,y)$ and $u_{2}=v_{2}(5.0,x,y)$. We extracted each function instance from a $64\times 64$ sub-mesh from the numerical solutions.
A.4 Torus Fluid
@@
where ${\bf x}\in[0,1]^{2}$, $t\in[0,10]$, $\nu=0.001$, ${\bf u}$ is the velocity field that $\nabla\cdot{\bf u}=0$, $w({\bf x},t)$ is the vorticity function, and $f({\bf x})$ is the source function that represents the external forces. We are interested in seven spatial functions: the initial condition $w_{0}$, the source function $f$, and the vortocity fields at time steps $t=2,4,6,8,10$, denoted as $w_{1}$, $w_{2}$, $w_{3}$, $w_{4}$, and $w_{5}$. We employed a parametric form for $w_{0}$ and $f$: $w_{0}({\bf x})=\left[\sin(\alpha_{1}\pi(x_{1}+\beta_{1})),\sin(\alpha_{2}\pi(x_{1}+\beta_{2}))\right]\cdot\mathbf{\Lambda}\cdot\left[\cos(\alpha_{1}\pi(x_{2}+\beta_{1})),\cos(\alpha_{2}\pi(x_{2}+\beta_{2}))\right]^{T}$, where $\alpha_{1},\alpha_{2}\sim\text{Uniform}(0.5,1)$, $\beta_{1},\beta_{2}\sim\text{Uniform}(0,1)$, and each $\left[\mathbf{\Lambda}\right]_{ij}\sim\text{Uniform}(-1,1)$, and $f(x_{1},x_{2})=0.1\left[a\sin(2\pi(x_{1}+x_{2}+c))+b\cos(2\pi(x_{1}+x_{2}+d)\right]$ where $a,b\sim\text{Uniform}(0,2)$ and $c,d\sim\text{Uniform}(0,0.5)$. We used the finite-difference solver as provided in (Li et al., 2020b) to obtain the solution of $\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. The spatial domain was discretized to a $128\times 128$ mesh. Again, to prepare the training and test datasets, we extracted the function instances from a $64\times 64$ sub-mesh.
Appendix B Hyperparameter Selection